a cell has an e.m.f. of 3v. when it is connected across a resistor of resistance 4ohms, a current of 0.5a passes through the current. calculator the internal resistance


1
Expert's answer
2022-02-10T15:13:36-0500

Gives "E=3V\\R=4\Omega\\i=0.5A"

We know that

"I=\frac{E}{R+r}"

"0.5=\frac{3}{4+r}"

"2+0.5r=3"

"r=\frac{10}{5}=2\Omega"